4 2/3 + 1 1/4 + 1 1/3 + 3/4 =
Found 2 solutions by fractalier, stanbon:
Answer by fractalier(6550)

To add
4 2/3 + 1 1/4 + 1 1/3 + 3/4 =
we would be best served to change all of these into a common denominator...12 in this case...we get
4 8/12 + 1 3/12 + 1 4/12 + 9/12 = 6 24/12 = 8

Answer by stanbon(75887)

4 2/3 + 1 1/4 + 1 1/3 + 3/4
-------------------
= 14/3 + 5/4 + 4/3 + 3/4
----------------------
lcd = 12
----
= (4*14 + 3*5 + 4*4 + 3*3)/ 12
-------
= (18*4 + 8*3)/12
------
= 96/12
----
= 8
------------
Cheers,
Stan H.
------------